OpenStudy (anonymous):

Can someone double check this for me? (x-10)(x+10) x * x = x^2 x * 10 = 10x -10 * x = -10x -10 * 10 = -100 = x^2 + 10x - 10x - 100 = x^2 + x - 100

OpenStudy (inkyvoyd):

No.

OpenStudy (anonymous):

10x-10x=0

OpenStudy (inkyvoyd):

= x^2 + 10x - 10x - 100 =x^2-100 in general, (a+b)(a-b)=a^2-b^2

Parth (parthkohli):

=> x(x+10) - 10(x + 10) => x^2 + 10 - 10x - 100 => x^2 - 10x - 90 Wrong!

OpenStudy (anonymous):

I'm using the FOIL method in this people..

Parth (parthkohli):

Answer would be the same, whatever method you use.